Prove for every integer n, n4 has the form 8m or 8m + 1 for some integer m. Hint: Use 2 cases.



1
Expert's answer
2022-03-18T09:18:16-0400

Let n is an even number


Then for some integer m, we can write


"n = 2m"


"n^{4}=(2m)^{4}"


"n^{4}=16m^{4}"


"n^{4}=8(2m^4)"


"n^{4}=8p" here "p=2m^4" an other integer


Let us consider n is an odd integer,


Then for some integer m, we can write,


"n = 2m +1"


Then


"n^{4} = (2m +1)^{4}"


Using binomial expansion,


"n^{4} = 16m^4+32m^3+24m^2+8m+1"


"n^{4} = 8(2m^4+4m^3+3m^2+m)+1"


"n^{4} = 8q+1" here "q=2m^4+4m^3+3m^2+m"


Hence proved that


For every integer "n", "n^4"  has the form "8m"  or "8m + 1" for some integer m.